# FSDAM: Few-Shot Driver Attention Modeling via Vision-Language Coupling

### v1

@@ 5.1 Baselines
We evaluate FSDAM on four driver attention datasets: BDD-A [10.1007/978-3-030-20873-8_42], DADA-2000 [fang2023dadadriverattentionprediction], DReyeVE [palazzi2018predicting], and W3D [zhou2025where], comparing against seven baselines: U²-Net [Qin_2020_PR], MINet [pang2020multiscaleinteractivenetworksalient], DBNet [9812524], DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic], and LLada [zhou2025where]. These models represent current approaches in saliency and driver attention research. All baselines were trained on the complete BDD-A training set and tested on each dataset’s official test split. For LLada [zhou2025where], we report published results LLada was trained on the W3D dataset.
For caption prediction, we evaluate FSDAM on the W3D benchmark and compare it with models representing the current spectrum of driver attention approaches, from fully supervised to zero-shot and few-shot methods. Fully trained baselines including GazeXplain [chen2024gazexplainlearningpredictnatural] and LLada [zhou2025where] utilize the complete W3D [zhou2025where] training dataset (70k samples). Zero-shot and in-context models such as Qwen-VL [bai2025qwen25vltechnicalreport] and LLaVA [10.5555/3666122.3667638] are evaluated without task-specific fine-tuning, with Qwen-VL [bai2025qwen25vltechnicalreport] using in-context learning via fine-tuning prompts. Few-shot baselines combine gaze prediction models with language models—DeepGazeIIE [linardos2021deepgazeiiecalibratedprediction] + LLaVA [10.5555/3666122.3667638] and MLNet [8356626] + LLaVA [10.5555/3666122.3667638] employ a two-stage approach where gaze heatmaps are first generated then fed to the language model, both trained on the same 90 BDD-A samples as our method. FSDAM (Ours) uses an end-to-end architecture trained on 90 BDD-A samples and is evaluated on the W3D test set across three driving scenarios: normal driving, safety-critical situations, and traffic accidents, demonstrating the effectiveness of our unified gaze-language approach against both data-rich and data-efficient alternatives.
@@ 5.2.1 Spatial Attention Prediction
Table 1 reports quantitative results for driver attention prediction across four datasets. All reproduced models were trained on BDD-A for consistency. FSDAM provides a strong balance between sample efficiency and predictive accuracy. With only 90 training samples, it reaches performance that aligns with or exceeds models trained on much larger datasets, including W3D with $\sim$70K frames, DADA-2000 with 658K frames, and DR(eye)VE with  [^5] K frames.
On BDD-A[palazzi2018predicting], FSDAM obtains the lowest KL divergence at 1.13. This improves over LLada[zhou2025where] by 2.6% and over DBNet[9812524] by 13%. It matches LLada[zhou2025where] in CC at 0.60 and ranks second in NSS at 4.10, indicating better spatial attention prediction. Compared to the fully supervised MINet[pang2020multiscaleinteractivenetworksalient], FSDAM improves CC by 30% and reduces KL divergence by 89%, which highlights its effectiveness with limited data.
Our cross-dataset evaluation demonstrates the strong transferability of our approach. On DR(eye)VE[palazzi2018predicting], FSDAM reaches the lowest KL divergence at 0.77, improving over LLada[zhou2025where] by 26% and DBNet[9812524] by 57%. It also achieves the highest SIM at 0.54, a 20% gain over U²-Net[Qin_2020_PR]. This shows that vision language alignment helps capture domain stable attention cues that transfer to highway scenarios.
@@
These results show that extensive supervision is not required for strong driver attention modeling. FSDAM ranks first in KL divergence on three of the four datasets with an average 15% margin over the next best method. This establishes the proposed few shot vision language framework as a practical and scalable solution for new driving domains where large gaze annotation sets are not available.
Table 1: Quantitative comparison of driver attention prediction across four datasets. All reproduced models were trained on BDD-A dataset. Bold indicates best performance, underlined indicates second best. * denotes results reported from original papers.
Method BDD-A DADA-2000 DR(eye)VE W3D
@@
U²-Net [Qin_2020_PR] 0.56 1.50 0.47 0.94 0.88 3.95 0.42 2.18 0.37 0.91 0.82 2.73 0.57 1.52 0.45 0.90 0.82 3.20 0.44 2.10 0.37 0.90 0.81 2.81
MINet [pang2020multiscaleinteractivenetworksalient] 0.46 10.3 0.16 0.89 0.82 3.67 0.32 10.45 0.14 0.82 0.73 2.02 0.44 8.87 0.36 0.84 0.80 2.65 0.35 10.51 0.14 0.83 0.77 2.54
DBNet [9812524] 0.57 1.30 0.41 0.95 0.91 4.41 0.41 1.89 0.29 0.93 0.85 3.08 0.48 1.79 0.29 0.91 0.85 3.71 0.47 1.77 0.33 0.93 0.86 3.50
DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] 0.47 9.62 0.21 0.97 0.75 2.56 0.42 10.26 0.19 0.96 0.73 2.23 0.67 8.78 0.47 0.91 0.84 2.74 0.53 9.70 0.32 0.93 0.81 2.35
LLada* [zhou2025where] 0.60 1.16 0.47 0.48 1.82 0.36 0.67 1.04 0.52 1.46 0.37 0.95 0.90 4.62
FSDAM (Ours) 0.60 1.13 0.43 0.96 0.91 4.10 0.44 1.64 0.33 0.92 0.86 2.95 0.61 0.77 0.54 0.91 0.84 4.04 0.53 1.27 0.44 0.91 0.87 3.50
@@ 5.2.2 Caption Prediction Analysis
Table 2 presents caption generation performance across three driving categories on the W3D dataset. Despite using only 90 training samples from BDD-A, our FSDAM achieves remarkable results that rival or exceed fully-trained baselines using the complete W3D dataset. In normal driving scenarios, FSDAM attains strong performance (BLEU: 0.42, METEOR: 0.37, CIDEr-R: 0.83) competitive with LLada (BLEU: 0.44, METEOR: 0.36, CIDEr-R: 0.96), which is trained on the full dataset. Notably, FSDAM demonstrates robust generalization to challenging scenarios, achieving competitive scores in safety-critical situations (BLEU: 0.35, METEOR: 0.33) and traffic accidents (BLEU: 0.33, METEOR: 0.35, CIDEr-R: 0.84). Compared to other few-shot baselines combining existing gaze models with LLaVA, FSDAM shows substantial improvements—achieving up to 3.2 times higher BLEU scores (0.42 vs. 0.13 for MLNet[pang2020multiscaleinteractivenetworksalient]+LLaVA in normal driving). Zero-shot approaches yield limited performance (BLEU << 0.21), underscoring the value of our few-shot learning framework, which effectively bridges the gap between zero-shot and full-data training regimes while preserving high performance efficiency.
@@ 5.3 Qualitative Results
Figure 4 presents qualitative comparisons against fully-supervised baselines U2-Net [Qin_2020_PR] and DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic]. Row (a) demonstrates peripheral hazard detection where FSDAM produces a broader attention distribution extending toward the right-side pedestrian area, while baselines concentrate more tightly on the stop sign region. In the complex intersection scenario (row c) with multiple pedestrians and vehicles, FSDAM generates the most spatially extensive attention coverage across agents, while U2-Net produces more concentrated predictions and DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] exhibits multiple distinct hot spots, potentially missing critical areas. Row (e) illustrates left-turn navigation where FSDAM effectively captures distributed attention spanning both the lead vehicle and the turning trajectory, demonstrating the benefit of vision-language alignment in understanding maneuver semantics, while baselines show more localized attention patterns. Row (d) shows all methods successfully localizing attention to the lane-changing vehicle’s brake light area, with comparable spatial precision across approaches.
Row (b) presents a challenging scenario where ground truth exhibits distributed attention across both the lead vehicle and a right-side pedestrian. All methods concentrate primarily on the vehicle, missing the distributed pattern. Despite this shared challenge, FSDAM demonstrates strong performance across diverse driving scenarios, producing attention predictions with superior spatial coverage in multi-agent situations (rows a, c, e) and comparable precision in focused attention tasks (row d). These results validate that our vision-language alignment approach enables effective few-shot learning, achieving spatial and semantic coherence competitive with fully-supervised methods trained on substantially larger datasets
Figure 4: Qualitative comparison of driver attention prediction on BDD-A test scenes showing the input image and attention heatmaps from ground truth, FSDAM trained on 90 samples, U2-Net[Qin_2020_PR], and DeepLabV3[chen2017rethinkingatrousconvolutionsemantic].
Figure 5: Driving scene and model output showing the input frame, its ground truth gaze map, and the corresponding FSDAM prediction. The model produces attention heatmaps over key regions and generates structured attention reasoning.
@@ 5.4 Few-Shot Learning Analysis
Figure 1 presents the impact of varying support set sizes on gaze prediction performance across six metrics. FSDAM demonstrates efficient learning dynamics, with substantial improvements from 1-shot to 5-shot settings. Performance gains are most pronounced between 1-shot and 3-shot (e.g., NSS increases from 2.32 to 3.64, CC improves from 0.41 to 0.53), after which the improvements plateau. Notably, our 5-shot model achieves performance (CC: 0.58, SIM: 0.43, NSS: 4.09, KL: 1.17) that closely approaches the full-data baseline (CC: 0.60, SIM: 0.43, NSS: 4.10, KL: 1.13), recovering 96.7% of full-data CC performance and 99.8% of NSS performance with only five support samples. The KL divergence shows consistent reduction from 1.74 to 1.17 as support size increases, indicating improved distribution matching. This analysis validates FSDAM’s data efficiency: meaningful gaze predictions emerge from minimal supervision, with performance rapidly saturating after 3-5 examples, making our approach practical for real-world deployment where annotated driver attention data is scarce.
@@ 5.5 Ablations
We conduct systematic ablation experiments to assess each component’s contribution in FSDAM. For gaze prediction, we evaluate on the official BDD-A test set. For caption generation, we evaluate on 49 curated samples with structured explanations. We compare four variants: Gaze-Only trains only gaze prediction; Caption-Only trains only caption generation; Shared Cross-Attention uses a single cross-attention module for both tasks; and Full FSDAM uses task-specific cross-attention modules with vision–language alignment.
Figure 6 shows a consistent performance hierarchy across both tasks. The shared cross-attention variant provides marginal gains over single-task baselines but degrades gaze prediction (4% higher KL divergence), indicating negative transfer. In contrast, full FSDAM achieves substantial improvements: 31.9% KL reduction, 43.3% SIM improvement, and 66.7% NSS improvement for gaze prediction, alongside 32.4% ROUGE-L, 7.1% CIDEr-R, and 3.4% BERTScore improvements for caption generation, all relative to single-task baselines. This hierarchy (single-task << shared << dual-pathway) validates three design principles: joint training enables mutual benefits through shared visual representations, task-specific cross-attention modules prevent negative transfer, and dual pathways with vision–language alignment enable effective few-shot learning.
Figure 6: Ablation study comparing four architectural variants. The dual-pathway configuration in FSDAM achieves the largest percentage gains for both gaze prediction and caption generation.

### v2

@@ 5.1 Baselines
We evaluate gaze prediction on four datasets: BDD-A [10.1007/978-3-030-20873-8_42], DADA-2000 [fang2023dadadriverattentionprediction], DR(eye)VE [palazzi2018predicting], and W3D [zhou2025where]. We compare against classical saliency/gaze baselines (U2-Net [Qin_2020_PR], MINet [pang2020multiscaleinteractivenetworksalient], DBNet [9812524], DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic]) and a joint gaze–language baseline (LLada [zhou2025where]). Unless noted otherwise, reproduced baselines are trained on the full BDD-A training split using the official implementations and evaluated on each dataset’s official test split. To provide a fair few-shot reference across both classical saliency and joint gaze–language baselines, we retrain U2-Net and LLada on the same 90-sample BDD-A subset as FSDAM (Table 1, Few-shot), ensuring differences are not attributable solely to training-data scale.
Table 1: In-domain driver attention prediction on BDD-A. Bold = best, underline = second best. results from original paper (trained on BDD-A). fine-tuned on 90 BDD-A samples.
[…]
Fully Supervised (Full BDD-A training set)
U²-Net [Qin_2020_PR] Full BDD-A 0.56 1.50 0.47 0.94 0.88 3.95
MINet [pang2020multiscaleinteractivenetworksalient] Full BDD-A 0.46 10.30 0.16 0.89 0.82 3.67
DBNet [9812524] Full BDD-A 0.57 1.30 0.41 0.95 0.91 4.41
DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] Full BDD-A 0.47 9.62 0.21 0.97 0.75 2.56
LLada [zhou2025where] Full BDD-A 0.60 1.16 0.47
Few-Shot (90 BDD-A samples)
U²-Net [Qin_2020_PR] BDD-A-90 0.10 3.43 0.13 0.74 0.68 0.63
LLada [zhou2025where] BDD-A-90 0.37 1.86 0.32 0.91 0.81 2.91
FSDAM (Ours) BDD-A-90 0.60 1.13 0.43 0.96 0.91 4.10
For captioning, we compare on W3D against fully supervised baselines (GazeXplain [chen2024gazexplainlearningpredictnatural], LLada [zhou2025where]; $\sim$}70k samples), zero-shot and in-context models (Qwen-VL [bai2025qwen25vltechnicalreport], LLaVA [10.5555/3666122.3667638]), and few-shot two-stage baselines pairing gaze models with LLaVA (DeepGazeIIE [linardos2021deepgazeiiecalibratedprediction]+LLaVA, MLNet [8356626]+LLaVA), trained on our same 90 BDD-A samples. Additional analyses and experiments are provided in the supplementary material.
@@ 5.2.1 Spatial Attention Prediction
Table 1 and Table 2 report quantitative results for driver attention prediction across four datasets. Table 1 evaluates in-domain performance on BDD-A under both fully supervised and few-shot regimes. Table 2 evaluates zero-shot transfer to DADA-2000, DR(eye)VE, and W3D, where all baselines are trained on full BDD-A and FSDAM uses only 90 samples. All reproduced models were trained on BDD-A for consistency. FSDAM provides a strong balance between sample efficiency and predictive accuracy. With only 90 training samples, it reaches performance that aligns with or exceeds models trained on much larger datasets, including W3D with $\sim$70K frames, DADA-2000 with 658K frames, and DR(eye)VE with  [^5] K frames.
In-domain Table 1. On BDD-A[10.1007/978-3-030-20873-8_42], FSDAM obtains the lowest KL divergence at 1.13. This improves over LLada[zhou2025where] by 2.6% and over DBNet[9812524] by 13%. It matches LLada[zhou2025where] in CC at 0.60 and ranks second in NSS at 4.10, indicating better spatial attention prediction. Compared to the fully supervised MINet[pang2020multiscaleinteractivenetworksalient], FSDAM improves CC by 30% and reduces KL divergence by 89%, which highlights its effectiveness with limited data. In the few-shot setting, classical saliency models such as U²-Net [Qin_2020_PR] degrade severely to CC=0.10, confirming that appearance-based features cannot generalize under extreme data scarcity. LLada [zhou2025where], despite sharing the same VLM backbone, also drops substantially to CC=0.37 and KL=1.86, demonstrating that vision-language alignment alone is insufficient without our dual-pathway few-shot design.
Cross-dataset transfer Table 2. Our cross-dataset evaluation demonstrates the strong transferability of our approach. On DR(eye)VE[palazzi2018predicting], FSDAM reaches the lowest KL divergence at 0.77, improving over LLada[zhou2025where] by 26% and DBNet[9812524] by 57%. It also achieves the highest SIM at 0.54, a 20% gain over U²-Net[Qin_2020_PR]. This shows that vision-language alignment helps capture domain stable attention cues that transfer to highway scenarios.
@@
Together, these results reveal that strong driver attention modeling does not require exhaustive gaze supervision. FSDAM consistently matches or outperforms fully supervised baselines across diverse driving scenarios, indicating that our dual-pathway vision-language design captures transferrable, semantically grounded attention cues absent from data-intensive saliency models.
Table 2: Zero-shot cross-dataset transfer for driver attention prediction. All baselines are trained on the full BDD-A training set; FSDAM uses only 90 BDD-A samples. No target-domain fine-tuning is applied for any method. Bold = best, underline = second best.
Method Training Data DADA-2000 DR(eye)VE W3D
@@
U²-Net [Qin_2020_PR] Full BDD-A 0.42 2.18 0.37 0.91 0.82 2.73 0.57 1.52 0.45 0.90 0.82 3.20 0.44 2.10 0.37 0.90 0.81 2.81
MINet [pang2020multiscaleinteractivenetworksalient] Full BDD-A 0.32 10.45 0.14 0.82 0.73 2.02 0.44 8.87 0.36 0.84 0.80 2.65 0.35 10.51 0.14 0.83 0.77 2.54
DBNet [9812524] Full BDD-A 0.41 1.89 0.29 0.93 0.85 3.08 0.48 1.79 0.29 0.91 0.85 3.71 0.47 1.77 0.33 0.93 0.86 3.50
DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] Full BDD-A 0.42 10.26 0.19 0.96 0.73 2.23 0.67 8.78 0.47 0.91 0.84 2.74 0.53 9.70 0.32 0.95 0.81 2.35
FSDAM (Ours) BDD-A-90 (few-shot) 0.44 1.64 0.33 0.92 0.86 2.95 0.61 0.77 0.54 0.91 0.84 4.04 0.53 1.27 0.44 0.91 0.87 3.50
@@ 5.2.2 Caption Prediction Analysis
Table 5 presents caption generation performance across three driving categories on the W3D dataset. Despite training under our few-shot regime, FSDAM achieves results approaching fully trained baselines trained on the full W3D dataset. In normal driving scenarios, FSDAM attains BLEU 0.42, METEOR 0.37, and CIDEr-R 0.83, closely approaching LLada (BLEU 0.44, METEOR 0.36, CIDEr-R 0.96) trained on $\sim$70k samples. Generalization holds across challenging scenarios, with competitive scores in safety-critical situations (BLEU: 0.35, METEOR: 0.33, CIDEr-R: 0.47) and traffic accidents (BLEU: 0.33, METEOR: 0.35, CIDEr-R: 0.84). Among few-shot baselines, two-stage approaches (DeepGazeIIE [linardos2021deepgazeiiecalibratedprediction]+LLaVA, MLNet [8356626]+LLaVA) achieve BLEU scores no higher than 0.26 across all scenarios, while FSDAM consistently exceeds 0.33 across normal driving, safety-critical, and accident scenarios. Zero-shot and in-context models remain limited across all scenarios (BLEU < 0.21), confirming that task-specific grounding is necessary for reliable driver caption generation.
@@ 5.3 Qualitative Results
Figure 4 compares FSDAM against fully-supervised U2-Net [Qin_2020_PR] and DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic]. In multi-agent scenarios (rows a, c, e), FSDAM produces broader attention coverage spanning peripheral pedestrians, intersection agents, and turning trajectories, whereas baselines concentrate on single targets or exhibit fragmented hot spots. In focused tasks (row d), all methods comparably localize attention to the lane-changing vehicle’s brake lights. Row (b) reveals a shared limitation: all methods miss the distributed ground-truth pattern across both the lead vehicle and a right-side pedestrian. Overall, FSDAM achieves superior spatial coverage in complex scenes while maintaining competitive precision, validating that vision-language alignment enables effective few-shot attention learning.
Figure 4: Qualitative comparison of driver attention prediction on BDD-A test scenes showing input image (left) and attention heatmaps from ground truth, FSDAM (90 samples), U2-Net[Qin_2020_PR], and DeepLabV3[chen2017rethinkingatrousconvolutionsemantic].
Figure 5: Few-shot learning performance on BDD-A. Metrics $C$, SIM, NSS, AUC-J, AUC-B, and KL (inverted) are min-max normalized to [0,1]; larger area = better.
Figure 6: Ablation study comparing four architectural variants. Dual-pathway FSDAM achieves the largest gains for both gaze prediction and caption generation.
@@ 5.4 Few-Shot Learning Analysis
Figure 6 presents the impact of varying support set sizes on gaze prediction performance across six metrics. FSDAM demonstrates efficient learning dynamics, with substantial improvements from 1-shot to 5-shot settings. Performance gains are most pronounced between 1-shot and 3-shot (e.g., NSS increases from 2.32 to 3.64, CC improves from 0.41 to 0.53), after which improvements plateau. Notably, our 5-shot model achieves performance (CC 0.58, SIM 0.43, NSS 4.09, KL 1.17) that closely approaches the full-data baseline (CC 0.60, SIM 0.43, NSS 4.10, KL 1.13), recovering 96.7% of full-data CC performance and 99.8% of NSS performance with only five support samples. KL divergence shows a consistent reduction from 1.74 to 1.17 as support size increases, indicating improved distributional alignment between predicted and ground-truth gaze. Importantly, the fact that SIM saturates early (SIM 0.43 at 5-shot matches the full-data SIM 0.43) suggests that the model rapidly learns the overall attention shape, while additional shots primarily refine correlation and peak alignment reflected by CC/NSS. This trend is consistent with our design goal: training-only vision–language alignment provides a strong semantic prior, and the gaze pathway then sharpens spatial localization as supervision increases. Overall, meaningful gaze predictions emerge from minimal supervision, with performance saturating after 3–5 examples.
@@ 5.5 Ablations
We conduct systematic ablation experiments to assess each component’s contribution in FSDAM. For gaze prediction, we evaluate on the official BDD-A test set. For caption generation, we evaluate on 49 curated samples with structured explanations, comprising all available test samples with complete four-component annotation. We compare four variants: Gaze-Only trains only gaze prediction; Caption-Only trains only caption generation; Shared Cross-Attention uses a single cross-attention module for both tasks; and Full FSDAM uses task-specific cross-attention modules with vision–language alignment.
Figure 6 shows a consistent performance hierarchy across both tasks. The shared cross-attention variant provides marginal gains over single-task baselines but degrades gaze prediction (4% higher KL divergence), indicating negative transfer. In contrast, full FSDAM achieves substantial improvements: 31.9% KL reduction, 43.3% SIM improvement, and 66.7% NSS improvement for gaze prediction, alongside 32.4% ROUGE-L, 7.1% CIDEr-R, and 3.4% BERTScore improvements for caption generation, all relative to single-task baselines. This hierarchy (single-task << shared << dual-pathway) validates three design principles: (1) multi-task training can improve overall performance by leveraging shared visual representations; (2) forcing both tasks to share a single cross-attention module induces negative transfer (e.g., higher KL divergence); and (3) decoupled pathways with training-only vision–language alignment yield the strongest and most consistent gains across both tasks.